A raft is made of 11 logs lashed together. Each is 44 cm in diameter and has a length of 6.0 m .

How many people can the raft hold before they start getting their feet wet, assuming the average person has a mass of 69 kg ? Do not neglect the weight of the logs. Assume the specific gravity of wood is 0.60.
Answer #1

11 logs

and diameter is 44 cm = 44/2

= 22 cm

= 0.22 m

the volume of the logs are

11 x 3.14 x ( 0.22 )2 x 6

= 10.030 m3

the density of log = specific gravity x density of water

here say the no of people is " n "

volume of log = 10.030 x 1000 ( since water density is 1000 )

volume of log x density of log + 69 n

= 10.030 x 1000 x 0.6 + 69 n

= 6018 + 69 n

now

10.030 x 1000 = 6018 + 69 n

10030 - 6018 = 69 n

n = 10030 - 6018 / 69

n = 58.144

n = 58 peoples approximately.